OpenStudy (anonymous):

Find all numbers for which the rational expression is undefined. v^2+13/v^2-3v-10

OpenStudy (anonymous):

Set the denominator equal to zero. Solve for v. v^2-3v-10 = 0 (v + 2) (v - 5) = 0 (v + 2) = 0 ... so v = -2 (v - 5) = 0 ... so v = 5

OpenStudy (anonymous):

The rational expression is undefined when v equals -2 or 5.

OpenStudy (anonymous):

When v equals -2 or 5 the denominator equals zero. This makes the rational expression undefined.